1 Calculate the kinetic energy of

(a) a 1 kg trolley travelling at 2 m s1,

6) a 2 g (0.002 kg) bullet travelling at 400 m s ,

(c)a 50 kgminicartravelling at 72 km h


1
Expert's answer
2021-02-05T04:30:04-0500

(a)

"KE=\dfrac{1}{2}mv^2=\dfrac{1}{2}\cdot1\ kg\cdot(2\ \dfrac{m}{s})^2=2\ J."

(b)

"KE=\dfrac{1}{2}\cdot0.002\ kg\cdot(400\ \dfrac{m}{s})^2=160\ J."

(c)

"KE=\dfrac{1}{2}\cdot50\ kg\cdot(72\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s})^2=10000\ J."